Find the general solution of the system
2x + 4y + 6z = 0
4x + 5y + 6z = 3
7x + 8y + 9z = 6:
1
Expert's answer
2020-10-21T14:19:07-0400

"\displaystyle

2x + 4y + 6z = 0 \hspace{1cm} (1)\\

4x + 5y + 6z = 3\hspace{1cm} (2)\\

7x + 8y + 9z = 6\hspace{1cm} (3)\\


(2) - (1) \\


2x + y = 3\\



(2) \times 3\\

4x + 5y + 6z = 3\\

12x + 15y + 18 = 9\hspace{1cm} (4)\\

(3) \times 2\\


14x + 16y + 18z = 12\hspace{1cm} (5)\\



(5) - (4)\\


2x + y = 3\\



\textsf{Let}\, x = n\, \textsf{where}\, n \in \mathbb{R}\\

y = 3 - 2x\\

y = 3 - 2n\\



\textsf{Substituting the value of}\, x \, \textsf{and}\, y \, \textsf{in}\, (1)\\

\begin{aligned}
2n + 4(3 -2n) + 6z &= 0\\
2n + 12 - 8n + 6z &= 0\\
6z &= 6n - 12\\
z &= n - 2
\end{aligned}\\


\therefore\textsf{The general solution of the simultaneous}\\\textsf{linear equation is}\\


x = n, y = 3 - 2n, z = n - 2,\, \textsf{for all}\, n \in \mathbb{R}\\"